Calculating IS, y=c+i+g+nx for:

y=50+.75(y-200)+260-20r+200+200-50e

Step by step please through algebra, Thanks
Answer #1

IS y = c + i + g + nx

y = 50 + 0.75(y - 200) + 260 - 20r + 200 + 200 - 50e

y = 710 + 0.75(y - 200) - 20r - 50e

y = 710 + 0.75y - 150 - 20r - 50e

y = 560 + 0.75y - 20r - 50e

y - 0.75y = 560 - 20r - 50e

0.25y = 560 - 20r - 50e

y/4 = 560 - 20r - 50e ( 0.25 = 1/4)

y = 2240 - 80r - 200e  

so IS : y = 2240 - 80r - 200e